There were 770 tickets sold for a ballet performance at the CCP. The tickets cost P2,500 for semi-VIP seats and P5000 for VIP seats. A total of P2,800,000 was collected. How many of each kind of ticket were sold?
Please give me the representation of your problem thank you very much....

Answer by checkley77(12844)

2,500x+5,000(770-x)=2,800,000
2,500x+3,850,000-5,000x=2,800,000
-2,500x=2,800,000-3,850,000
-2,500x=-1,050,000
x=-1,050,000/-2,500
x=420 semi-VIP tickets sold.
770-420=350 VIP tickets sold.
Proof:
2,500*420+5,000*350=2,800,000
1,050,000+1,750,000=2,800,000
2,800,000=2,800,000